Fermi National Laboratory
 

Accelerator Update for November 21, 2007 - November 26, 2007

 

Wednesday November 21

The day shift began with experts conducting Tevatron (TeV) studies, with Operations monitoring the Recycler stash (270.6E10), with the Antiproton Source (Pbar) stacking (26.4E10), with the Main Injector (MI), Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for beam, with NuMI off due to an absorber water skid controls problem, and with the SciBooNE / MiniBooNE experiments taking beam.

Experts completed their NuMI repairs. Operations resumed sending beam to NuMI at 9:17 AM.

At 10:10 AM, the Booster East Anode Power Supply (EAPS) tripped off. Later, an expert reported that a Booster RF station had tripped (BRF2) and caused the EAPS trip. He bypassed the station and brought BRF11 on line.

Operations put the TeV through a dry squeeze at 10:58 AM.

Operations put the TeV into shot setup at 1:06 PM.

Operations reported at 4:02 PM that the Recycler had stashed antiprotons once during the day shift.

Operations established store 5741 at 4:11 PM with an initial luminosity of 178.16E30.

Experts reported at 5:02 PM that they had brought the Booster RF system back online and completed work on a MI-60 power supply. Sci/MiniBooNE resumed taking beam.

At 8:24 PM, the Pbar beam permit tripped off due to an AP0 target hall ventilation alarm. Experts investigated and reported that there was nothing wrong. Pbar resumed stacking at 9:52 PM.

Operations reported at 11:54 PM that the Recycler had stashed three times during the evening shift.

Thursday November 22

The midnight shift began with Operations monitoring store 5741 (70.90E30) and the stash (176.8E10), with Pbar stacking (40.6E10), and with NuMI and Sci /MiniBooNE taking beam.

Operations reported at 8 AM that the Recycler had stashed three times during the midnight shift.

Operations terminated store 5741 at 11:04 AM.

Operations put the TeV into shot setup at 1:04 PM.

Operations established store 5743 at 3:37 PM with an initial luminosity of 168.16E30.

Operations reported at 3:56 PM that the Recycler had stashed three times during the day shift.

At 7:05 PM, LRF3 tripped off for the first of two times in seven minutes with no indication of a fault. It reset.

At 7:11 PM, the NuMI beam permit tripped off due to a power supply trip (E:V118). This was the first of five trips in about five hours. It reset each time.

Operations reported at 11:59 PM that the Recycler had stashed three times during the evening shift.

Friday November 23

The midnight shift began with Operations monitoring store 5743 (63.87E30) and a stash (190.8E10), with Pbar stacking (12.2E10), and with NuMI and Sci/MiniBooNE taking beam.

At 1:16 AM, E:V118 caused NuMI to trip off for the first of eight times over the next five hours. It reset each time.

Operations terminated store 5743 at 5:32 AM.

Operations prepared CDF and the TeV for a planned access for maintenance and repairs at 5:50 AM.

D0 went in on access at 7:06 AM.

At 7:59 AM, a cryo system expert reported that he found a large ice ball on some TeV sector A4 leads and broken hoses on the lower connections. They began warming up the area to make repairs.

Operations reported at 8:07 AM that the Recycler had stashed two times during the midnight shift.

At 8:17 AM, Operations reported that E:V118 had tripped off for the first of four times in about 45 minutes. It reset. A NuMI expert came in at 10:31 AM to investigate.

At 10:35 AM, Operations reported that a NuMI expert came in to investigate evaporator problems.

D0 came out of access at 11:04 AM.

A cryo system expert reported at 11:17 AM that the TeV sector A4 work was done.

At 12:12 PM, LRF3 tripped off. It reset.

At 1:26 PM, E:V118 began tripping off again. Operations called in a power supply expert.

At 1:29 PM, Operations reported problems with the TeV sector C3 wet engine. An expert said that a fan for its controller had failed, but since it would take about four hours to replace and the controller wasn’t overheating, he would mark it for repair at a later time.

Operations reported at 4:12 PM that the Recycler had stashed once during the day shift.

At 5:07 PM, LRF3 tripped off. It reset.

At 6:33 PM, LRF3 tripped off for the first of four times in about 25 minutes. It reset.

CDF came out of access at 6:54 PM.

Operations put the TeV through a dry squeeze at 8:32 PM.

Operations put the TeV through a wet squeeze at 10:46 PM.

Operations put the TeV into shot setup at 11:25 PM.

At 11:51 PM, LRF3 tripped off. It reset.

Saturday November 24

The midnight shift began with the TeV in shot setup, with Operations monitoring a stash (299.5E10), with Pbar stacking (193.5E10), and with NuMI and Sci/MiniBooNE taking beam.

Operations established store 5745 at 1:58 AM with an initial luminosity of 161.4E30.

At 3:33 AM, TeV store 5745 aborted. Operators contacted experts and later called in two power supply experts.

Operations reported at 7:58 AM that the Recycler had stashed three times during the midnight shift.

TeV experts reported at 9:50 AM that a coordinating device at TeV sector B3 registered a VCB permit trip, which was caused by power failure on the SPU at the bottom of the same equipment rack. The power supply experts swapped the coordinating device and the SPU with other units.

Operations put the TeV into shot setup at 12:08 PM.

At 12:43 PM, LRF3 tripped off for the first of three times in six minutes. It reset.

At 2:08 PM, Operations reported that the TeV suffered a sector A1 quench. The D0 detector had pulled an abort due to bad beam positions. TeV experts came in to investigate.

At 4 PM, LRF3 tripped off for the first of three times in 50 minutes. It reset.

Operations put the TeV through a dry squeeze at 7:25 PM.

Operations reported at 7:59 PM that the Recycler had stashed three times during the last 12 hours.

Operations put the TeV into shot setup at 10:05 PM.

Sunday November 25

The midnight shift began with the TeV in shot setup, with Operations monitoring a stash, with Pbar stacking, and with NuMI and Sci/MiniBooNE taking beam.

Operations established store 5748 at 12:05 AM with an initial luminosity of 108.5E30.

At 12:22 AM, LRF3 tripped off. It reset.

Operations reported at 8:03 AM that the Recycler had stashed five times during the last 12 hours.

At 8:18 AM, LRF3 tripped off. It reset.

At 9:17 AM, E:V118 tripped off. It reset.

At 10:08 AM, LRF3 tripped off for the first of three times in about an hour. It reset.

Operations terminated store 5748 at 12:01 PM.

Operations put the TeV into shot setup at 12:11 PM.

Operations established store 5749 at 2:07 PM with an initial luminosity of 145.17E30.

At 2:14 PM, LRF3 tripped off for the first of six times in about 4.5 hours. It reset.

Operations reported at 8 PM that the Recycler had stashed four times during the last 12 hours.

At 8:55 PM, Operations reported that they had to add some pressure to LRF3’s LCW water skid.

At 10:14 PM, E:V118 tripped off. It reset.

Monday November 26

The midnight shift began with Operations monitoring store 5748 and a stash, with Pbar stacking, and with NuMI and Sci/MiniBooNE taking beam.

At 1:09 AM, LRF3 tripped off. It reset.

Operations terminated store 5748 at 3:59 AM.

At 5:23 AM, E:V118 tripped off. It reset.

Operations put the TeV into shot setup at 4:11 AM.

Operations established store 5750 at 6:12 AM with an initial luminosity of 153.1E30.

At 7:03 AM, E:V118 tripped off. It reset.

Operations reported at 8:05 AM that the Recycler had stashed four times during the last 12 hours.

The Plans for Monday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/26/2007 -->    email Fermilab